OpenStudy (anonymous):

25 square 4 reduce the index as low as possible.

OpenStudy (anonymous):

\[\sqrt[4]{25?}\]

OpenStudy (anonymous):

\[\sqrt[4]{5^2}=\sqrt{5}\]

OpenStudy (anonymous):

i.e cancel the 2 in the index and radicand